# Volume-law entanglement fragmentation of quasiparticles

### v1

@@ Abstract
We study the entanglement entropy in quasiparticle states where certain unit patterns are excited repeatedly and sequentially in momentum space. We find that in the scaling limit, each unit pattern contributes independently and universally to the entanglement, leading to a volume-law scaling of the entanglement entropy. This characteristic of volume-law entanglement fragmentation is numerically confirmed in both fermionic and bosonic chains. We derive an analytical formula for fermions, which can also be applied to the spin-1/2 XXZ chain with appropriate identifications.
I Introduction
@@
Quantum entanglement has been instrumental in various areas of physics, such as quantum information theory, condensed matter physics, and high-energy physics Amico:2007ag ; Eisert:2008ur ; Calabrese:2009bph ; Laflorencie:2015eck ; Rangamani:2016dms ; Witten:2018lha . Notably, entanglement adheres to distinct scaling laws in various phases of quantum systems, which allows it to serve as an indicator of quantum phase transitions Osterloh:2002auc ; Osborne:2002zz . In the context of one-dimensional quantum spin chains with energy gaps, the ground state entanglement entropy typically conforms to the area law. Conversely, in critical spin chains, the ground state entropy adheres to a logarithmic law, with the proportionality constant being related to the central charge Holzhey:1994we ; Vidal:2002rm ; Latorre:2003kg ; Korepin:2004zz ; Calabrese:2004eu .
@@
The study of the time evolution of entanglement entropy following a quantum quench in many-body systems Calabrese:2005in has motivated research into entanglement entropy in excited states. This includes both low-energy states Alba:2009th ; Alcaraz:2011tn ; Berganza:2011mh ; Pizorn:2012aut ; Essler:2012rai ; Berkovits:2013mii ; Storms:2013wzf ; Calabrese:2014ntv ; Molter:2014qsb ; Taddia:2016dbm ; Castro-Alvaredo:2018dja ; Castro-Alvaredo:2018bij ; Castro-Alvaredo:2019irt ; Castro-Alvaredo:2019lmj ; Jafarizadeh:2019xxc ; You:2020osa ; Zhang:2020vtc ; Zhang:2020dtd ; Zhang:2021bmy , where the entropy adheres to either the area law or logarithmic law, and high-energy states Deutsch:2012vkm ; Santos:2012rss ; Beugeling:2015plv ; Garrison:2015lva ; Nakagawa:2017yiw ; Lu:2017tbo ; Miao:2019xpp , where the entropy typically aligns with the volume law. In Vidmar:2017uux ; Vidmar:2017pak ; Vidmar:2018rqk ; Huang:2019dxk ; Lydzba:2020qfx ; Bianchi:2021aui , the average entanglement entropy across the spectrum has been employed to characterize the universal properties of chaotic and integrable systems.
@@
Entanglement of excited states in integrable models can often be understood in terms of quasiparticles Calabrese:2005in ; Alba:2017hlc ; Alba:2017lvc . When a finite number of quasiparticles with large energy and significant momentum differences are excited, a semiclassical picture can describe the entanglement entropy Castro-Alvaredo:2018dja ; Castro-Alvaredo:2018bij ; Castro-Alvaredo:2019irt ; Castro-Alvaredo:2019lmj . For earlier related works, see Pizorn:2012aut ; Berkovits:2013mii ; Molter:2014qsb . On the other hand, when the momentum differences between the excited quasiparticles are small, there are strong coherence effects among the quasiparticles, which results in significant corrections to the semiclassical picture of the entanglement entropy Zhang:2020vtc ; Zhang:2020dtd ; Zhang:2021bmy .
@@
In this paper, we study the entanglement entropy in states with a large number of quasiparticles, where certain unit patterns are excited sequentially and repeatedly in the momentum space. We consider a subsystem consisting of a consecutive block within circular quantum systems, including both free and interacting fermionic and bosonic chains, as well as spin-1/2 XXZ chains. Generally, the entanglement entropy in such states follows a volume law. In Alba:2009th , analytical expressions were obtained for fermionic chains in cases where the subsystem size is much smaller than the size of the entire system. In the case of free fermionic chains, we obtain analytical expressions for subsystems of arbitrary sizes, which, in the scaling limit, also apply to interacting fermionic chains and spin-1/2 XXZ chains.
@@
II Free fermions
@@
$H=\sum_{j=1}^{L}\Big{(}a_{j}^{\dagger}a_{j}-\frac{1}{2}\Big{)},$ (1)
with the fermionic modes $a_{j}$ and $a_{j}^{\dagger}$ satisfying $\{a_{j_{1}},a_{j_{2}}\}=\{a^{\dagger}_{j_{1}},a^{\dagger}_{j_{2}}\}=0$ and $\{a_{j_{1}},a^{\dagger}_{j_{2}}\}=\delta_{j_{1}j_{2}}$. The excited states with translational invariance are generated by the global modes
$b_{k}^{\dagger}=\frac{1}{\sqrt{L}}\sum_{j=1}^{L}a_{j}^{\dagger}\mathrm{e}^{\frac{2\pi\mathrm{i}jk}{L}},~{}~{}b_{k}=\frac{1}{\sqrt{L}}\sum_{j=1}^{L}a_{j}\mathrm{e}^{-\frac{2\pi\mathrm{i}jk}{L}},$ (2)
@@
In the free fermionic chain, one has Cheong:2002ukf ; Vidal:2002rm ; Peschel:2002jhw ; Latorre:2003kg
@@
$S_{L,L_{A},K}^{\text{fer}}=\textrm{tr}_{A}[-C_{L,L_{A},K}\log C_{L,L_{A},K}$
$-(1-C_{L,L_{A},K})\log(1-C_{L,L_{A},K})],$ (4)
where the $L_{A}\times L_{A}$ matrix has entries
$[C_{L,L_{A},K}]_{j_{1}j_{2}}=\langle a_{j_{1}}^{\dagger}a_{j_{2}}\rangle_{K}=h_{L,j_{2}-j_{1},K},$ (5)
@@
$h_{L,j,K}=\frac{1}{L}\sum_{k\in K}\mathrm{e}^{\frac{2\pi\mathrm{i}jk}{L}}.$ (6)
@@
II.1 Fully occupied states
@@
From (6) we get the correlation function
$h_{pl,j,p\kappa}=\delta_{(j\,{\rm mod}\,p)=0}h_{l,j/p,\kappa},$ (9)
@@
We parameterize the subsystem length as $L_{A}=\alpha p+a$, with $\alpha=0,1,\cdots,l-1$ and $a=0,1,\cdots,p-1$, and the corresponding correlation matrix $C_{L,L_{A},K}$ is similar to
$C_{pl,\alpha p+a,p\kappa}\sim\underbrace{C_{l,\alpha+1,\kappa}\oplus\cdots}_{a}\oplus\underbrace{C_{l,\alpha,\kappa}\oplus\cdots}_{p-a}.$ (10)
@@
$\rho_{pl,\alpha p+a,p\kappa}\sim\underbrace{\rho_{l,\alpha+1,\kappa}\otimes\cdots}_{a}\otimes\underbrace{\rho_{l,\alpha,\kappa}\otimes\cdots}_{p-a}.$ (11)
There are $p$ copies of the unit pattern $\kappa$, and each copy is effectively confined in a circle of $l$ sites. The subsystem has $L_{A}=\alpha p+a$ sites, and they are effectively composed of subsystems of the $p$ copies of the small systems. An example scheme of the entanglement fragmentation is shown in Fig. 1.
@@
From the RDM (11), we get the decomposition of the Rényi entropy and entanglement entropy
$S_{pl,\alpha p+a,p\kappa}^{(n),{\textrm{fer}}}=aS_{l,\alpha+1,\kappa}^{(n),{\textrm{fer}}}+(p-a)S_{l,\alpha,\kappa}^{(n),{\textrm{fer}}}.$ (12)
This inspired us to define the contributions to the entanglement from each unit pattern as
$s^{(n),{\textrm{fer}}}_{l,\kappa}(x)\equiv\frac{S^{(n),{\textrm{fer}}}_{pl,xpl,p\kappa}}{p},$ (13)
@@
$s^{(n),{\textrm{fer}}}_{l,\kappa}(x)=yS^{(n),{\textrm{fer}}}_{l,\alpha+1,\kappa}+(1-y)S^{(n),{\textrm{fer}}}_{l,\alpha,\kappa},$ (14)
@@
[…]
Note that these formulas are exact, valid for any $p$, $l$, $\alpha$, $a$, and $\kappa$. Furthermore, the function $s_{l,\kappa}^{(n),{\textrm{fer}}}(x)$ is independent of $p$. For finite $l$ and large $p$, i.e. $l\ll p$, the Rényi entropy and entanglement entropy (12) follow the volume law. Generally, the formula $s_{l,\kappa}^{(n),{\textrm{fer}}}(x)$ is a piecewise function of $x\in(0,1)$ with $l$ pieces, each piece has length $\frac{1}{l}$, and within each piece the function is linear.
We use $|\kappa|$ to denote the number of excited quasiparticles in the set $\kappa$. In the special limit, $|\kappa|\ll l$, the formula (14) becomes
$\lim_{l\to+\infty}s^{(n),{\textrm{fer}}}_{l,\kappa}(x)=\lim_{l\to+\infty}S_{l,xl,\kappa}^{(n),{\textrm{fer}}}.$ (16)
In the same limit, the formula (13) becomes
$\lim_{l\to+\infty}s^{(n),{\textrm{fer}}}_{l,\kappa}(x)=\lim_{l\to+\infty}\frac{S_{pl,xpl,p\kappa}^{(n),{\textrm{fer}}}}{p}.$ (17)
The results of (16) and (17) should be the same and this is consistent with the fact, found in Zhang:2020vtc ; Zhang:2020dtd ; Zhang:2021bmy that quasiparticles with large momentum differences have independent contributions to the entanglement.
For $\kappa=\{0\}$, with $l=2,3,\cdots$, there are Pizorn:2012aut ; Berkovits:2013mii ; Molter:2014qsb ; Castro-Alvaredo:2018dja ; Castro-Alvaredo:2018bij
$S_{l,l_{A},\{0\}}^{(n),{\textrm{fer}}}=-\frac{1}{n-1}\log\Big{[}\Big{(}\frac{l_{A}}{l}\Big{)}^{n}+\Big{(}1-\frac{l_{A}}{l}\Big{)}^{n}\Big{]},$
$S_{l,l_{A},\{0\}}^{{\textrm{fer}}}=-\frac{l_{A}}{l}\log\frac{l_{A}}{l}-\Big{(}1-\frac{l_{A}}{l}\Big{)}\log\Big{(}1-\frac{l_{A}}{l}\Big{)},$ (18)
with which the formula (12) is the same as the comb entropy defined in Keating:2006psq after one takes the position-momentum duality Lee:2014nra ; Carrasco:2017eul . For $\kappa=\{k_{1},k_{2}\}$ with $l=3,4,\cdots$, the analytical formulas for $S_{l,l_{A},\kappa}^{(n),{\textrm{fer}}}$ can be found in Zhang:2020vtc ; Zhang:2020dtd ; Zhang:2021bmy . Examples of the expression (14) for the entanglement entropy from each unit pattern are shown in Fig. 2.
Figure 2: The formula (14) of the contribution to the entanglement entropy from each unit pattern in the momentum space of a free fermionic chain.
II.2 Partially occupied states
@@
In the scaling limit, we conjecture that each unit pattern still contributes independently to the entanglement entropy, and there is
$S^{(n),{\textrm{fer}}}_{L,xL,p\kappa}\approx ps^{(n),{\textrm{fer}}}_{l,\kappa}(x),$ (21)
with the universal function $s^{(n),{\textrm{fer}}}_{l,\kappa}(x)$ being defined in (13) and taking the form (14). We further get
$\lim_{L\to+\infty}\frac{S_{L,xL,p\kappa}^{(n),{\textrm{fer}}}}{L}=zs_{l,\kappa}^{(n),{\textrm{fer}}}(x),$ (22)
with the definition
$z\equiv\lim_{L\to+\infty}\frac{p}{L}\in\Big{[}0,\frac{1}{l}\Big{]}.$ (23)
It is easy to confirm the formula (22) numerically.
II.3 States with mixed occupancy

### v2

@@ Abstract
We investigate the entanglement entropy in quantum states featuring repeated sequential excitations of unit patterns in momentum space. In the scaling limit, each unit pattern contributes independently and universally to the entanglement entropy, leading to a characteristic volume-law scaling. Crucially, this universal contribution remains identical for both free and interacting models, enabling decomposition of the total entanglement into pattern-specific components. Numerical verification in fermionic and bosonic chains confirms this volume-law fragmentation phenomenon. For fermionic systems, we derive analytical expressions where many-body entanglement becomes expressible through few-body entanglement components. Notably, this analytical framework extends to spin-1/2 XXZ chains through appropriate identifications.
Center for Joint Quantum Studiesand Department of Physics, School of Science,
Tianjin University, 135 Yaguan Road, Tianjin 300350, China
1 Introduction
@@
Quantum entanglement has been instrumental in various areas of physics, such as quantum information theory, condensed matter physics, and high-energy physics [1, 2, 3, 4, 5, 6]. Notably, entanglement adheres to distinct scaling laws in various phases of quantum systems, which allows it to serve as an indicator of quantum phase transitions [7, 8]. In the context of one-dimensional quantum spin chains with energy gaps, the ground state entanglement entropy typically conforms to the area law. Conversely, in critical spin chains, the ground state entropy adheres to a logarithmic law, with the proportionality constant being related to the central charge [9, 10, 11, 12, 13].
@@
The study of the time evolution of entanglement entropy following a quantum quench in many-body systems [14] has motivated research into entanglement entropy in excited states. This includes both low-energy states [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], where the entropy adheres to either the area law or logarithmic law, and high-energy states [34, 35, 36, 37, 38, 39, 40, 41], where the entropy typically aligns with the volume law. In [42, 43, 44, 45, 46, 47], the average entanglement entropy across the spectrum has been employed to characterize the universal properties of chaotic and integrable systems.
@@
Entanglement of excited states in integrable models can often be understood in terms of quasiparticles [14, 48, 49]. When a finite number of quasiparticles with large energy and significant momentum differences are excited, a semiclassical picture can describe the entanglement entropy [25, 26, 27, 28]. For earlier related works, see [18, 20, 23]. On the other hand, when the momentum differences between the excited quasiparticles are small, there are strong coherence effects among the quasiparticles, which results in significant corrections to the semiclassical picture of the entanglement entropy [31, 32, 33].
@@
In this paper, we study the entanglement entropy in states with a large number of quasiparticles, where certain unit patterns are excited sequentially and repeatedly in the momentum space. We consider a subsystem consisting of a consecutive block within circular quantum systems, including both free and interacting fermionic and bosonic chains, as well as spin-1/2 XXZ chains. Generally, the entanglement entropy in such states follows a volume law. In [15], analytical expressions were obtained for fermionic chains in cases where the subsystem size is much smaller than the size of the entire system. In the case of free fermionic chains, we obtain analytical expressions for subsystems of arbitrary sizes, which, in the scaling limit, also apply to interacting fermionic chains and spin-1/2 XXZ chains.
@@
In the remaining part of the paper, we consider free and interacting fermions in sections 2 and 3, spin-1/2 XXZ chain in section 4, and free and interacting bosons in sections 5 and 6. We conclude with discussions in section 7.
2 Free fermions
@@
$H=\sum_{j=1}^{L}\Big{(}a_{j}^{\dagger}a_{j}-\frac{1}{2}\Big{)},$ (2.1)
with the fermionic modes $a_{j}$ and $a_{j}^{\dagger}$ satisfying $\{a_{j_{1}},a_{j_{2}}\}=\{a^{\dagger}_{j_{1}},a^{\dagger}_{j_{2}}\}=0$ and $\{a_{j_{1}},a^{\dagger}_{j_{2}}\}=\delta_{j_{1}j_{2}}$. The Hamiltonian is already in diagonal form, so we can derive the following analytical expressions. The excited states with translational invariance are generated by the global modes
$b_{k}^{\dagger}=\frac{1}{\sqrt{L}}\sum_{j=1}^{L}a_{j}^{\dagger}\mathrm{e}^{\frac{2\pi\mathrm{i}jk}{L}},~~b_{k}=\frac{1}{\sqrt{L}}\sum_{j=1}^{L}a_{j}\mathrm{e}^{-\frac{2\pi\mathrm{i}jk}{L}},$ (2.2)
@@
In the free fermionic chain, one has [50, 10, 51, 11]
@@
$S_{L,L_{A},K}^{\text{fer}}=\textrm{tr}_{A}[-C_{L,L_{A},K}\log C_{L,L_{A},K}-(1-C_{L,L_{A},K})\log(1-C_{L,L_{A},K})],$ (2.4)
which applies to any Hamiltonian for which the number of excitations is conserved, with the Hamiltonian (2.1) as a specific example. The $L_{A}\times L_{A}$ matrix $C_{L,L_{A},K}$ has entries
$C_{L,L_{A},K}]_{j_{1}j_{2}}=\langle a_{j_{1}}^{\dagger}a_{j_{2}}\rangle_{K}=h_{L,j_{2}-j_{1},K},$ (2.5)
@@
$h_{L,j,K}=\frac{1}{L}\sum_{k\in K}\mathrm{e}^{\frac{2\pi\mathrm{i}jk}{L}}.$ (2.6)
@@
2.1 Fully occupied states
@@
From (2.6) we get the correlation function
$h_{pl,j,p\kappa}=\delta_{j\,{\rm mod}\,p,0}h_{l,j/p,\kappa},$ (2.9)
@@
We parameterize the subsystem length as $L_{A}=\alpha p+a$, with $\alpha=0,1,\cdots,l-1$ and $a=0,1,\cdots,p-1$, and the corresponding correlation matrix $C_{L,L_{A},K}$ can be transformed to the following form via similarity transformation
$C_{pl,\alpha p+a,p\kappa}\sim\underbrace{C_{l,\alpha+1,\kappa}\oplus\cdots\oplus C_{l,\alpha+1,\kappa}}_{a}\oplus\underbrace{C_{l,\alpha,\kappa}\oplus\cdots\oplus C_{l,\alpha,\kappa}}_{p-a}.$ (2.10)
@@
$\rho_{pl,\alpha p+a,p\kappa}\sim\underbrace{\rho_{l,\alpha+1,\kappa}\otimes\cdots\otimes\rho_{l,\alpha+1,\kappa}}_{a}\otimes\underbrace{\rho_{l,\alpha,\kappa}\otimes\cdots\otimes\rho_{l,\alpha,\kappa}}_{p-a}.$ (2.11)
There are $p$ copies of the unit pattern $\kappa$, and each copy is effectively confined in a circle of $l$ sites. The subsystem has $L_{A}=\alpha p+a$ sites, and it effectively decomposes into subsystems across $p$ copies of small systems. An example scheme of the entanglement fragmentation is shown in Fig. 1.
@@
From the RDM (2.11), we get the decomposition of the Rényi entropy and entanglement entropy
$S_{pl,\alpha p+a,p\kappa}^{(n),{\textrm{fer}}}=aS_{l,\alpha+1,\kappa}^{(n),{\textrm{fer}}}+(p-a)S_{l,\alpha,\kappa}^{(n),{\textrm{fer}}}.$ (2.12)
Note that we have used the integer $p$ as the number of pattern repetitions. This inspired us to define the contributions to the entanglement from each unit pattern as
$s^{(n),{\textrm{fer}}}_{l,\kappa}(x)\equiv\frac{S^{(n),{\textrm{fer}}}_{pl,xpl,p\kappa}}{p},$ (2.13)
@@
$s^{(n),{\textrm{fer}}}_{l,\kappa}(x)=yS^{(n),{\textrm{fer}}}_{l,\alpha+1,\kappa}+(1-y)S^{(n),{\textrm{fer}}}_{l,\alpha,\kappa},$ (2.14)
@@
$y\equiv(lx{\rm\,mod\,}1)\in[0,1).$ (2.15)
Here we use $\lfloor lx\rfloor$ to denote the floor function, which gives the largest integer that is smaller than or equal to the argument $lx$. Note that these formulas are exact, valid for any $p$, $l$, $\alpha$, $a$, and $\kappa$. Furthermore, the function $s_{l,\kappa}^{(n),{\textrm{fer}}}(x)$ is independent of $p$. For finite $l$ and large $p$, i.e. $l\ll p$, the Rényi entropy and entanglement entropy (2.12) follow the volume law. Generally, the formula $s_{l,\kappa}^{(n),{\textrm{fer}}}(x)$ is a piecewise function of $x\in(0,1)$ with $l$ pieces, each piece has length $\frac{1}{l}$, and within each piece the function is linear.
We use $|\kappa|$ to denote the number of excited quasiparticles in the set $\kappa$. In the special limit, $|\kappa|\ll l$, the formula (2.14) becomes
$\lim_{l\to+\infty}s^{(n),{\textrm{fer}}}_{l,\kappa}(x)=\lim_{l\to+\infty}S_{l,xl,\kappa}^{(n),{\textrm{fer}}}.$ (2.16)
In the same limit, the formula (2.13) becomes
$\lim_{l\to+\infty}s^{(n),{\textrm{fer}}}_{l,\kappa}(x)=\lim_{l\to+\infty}\frac{S_{pl,xpl,p\kappa}^{(n),{\textrm{fer}}}}{p}.$ (2.17)
The results of (2.16) and (2.17) should be the same and this is consistent with the fact, found in [31, 32, 33] that quasiparticles with large momentum differences have independent contributions to the entanglement.
For $\kappa=\{0\}$, with $l=2,3,\cdots$, there are [18, 20, 23, 25, 26]
$\hskip-8.53581ptS_{l,l_{A},\{0\}}^{(n),{\textrm{fer}}}=-\frac{1}{n-1}\log\Big{[}\Big{(}\frac{l_{A}}{l}\Big{)}^{n}+\Big{(}1-\frac{l_{A}}{l}\Big{)}^{n}\Big{]},$
$\hskip-8.53581ptS_{l,l_{A},\{0\}}^{{\textrm{fer}}}=-\frac{l_{A}}{l}\log\frac{l_{A}}{l}-\Big{(}1-\frac{l_{A}}{l}\Big{)}\log\Big{(}1-\frac{l_{A}}{l}\Big{)},$ (2.18)
with which the formula (2.12) is the same as the comb entropy defined in [52] after one takes the position-momentum duality [53, 54]. For $\kappa=\{k_{1},k_{2}\}$ with $l=3,4,\cdots$, the analytical formulas for $S_{l,l_{A},\kappa}^{(n),{\textrm{fer}}}$ can be found in [31, 32, 33]. Examples of the expression (2.14) for the entanglement entropy from each unit pattern are shown in Fig. 2.
Figure 2: The formula (2.14) of the contribution to the entanglement entropy from each unit pattern in the momentum space of a free fermionic chain. In the left panel we use the unit pattern $\kappa=\{0\}$, and in the right panel we use the unit pattern $\kappa=\{0,1\}$. In each panel, the value of $l$ increases from lower to upper curves.
2.2 Partially occupied states
@@
In the scaling limit, we conjecture that each unit pattern continues to contribute independently to the entanglement entropy, and there is
$S^{(n),{\textrm{fer}}}_{L,xL,p\kappa}\approx ps^{(n),{\textrm{fer}}}_{l,\kappa}(x),$ (2.21)
with the universal function $s^{(n),{\textrm{fer}}}_{l,\kappa}(x)$ being defined in (2.13) and taking the form (2.14). We further get
$\lim_{L\to+\infty}\frac{S_{L,xL,p\kappa}^{(n),{\textrm{fer}}}}{L}=zs_{l,\kappa}^{(n),{\textrm{fer}}}(x),$ (2.22)
with the definition of the repetition ratio
$z\equiv\lim_{L\to+\infty}\frac{p}{L}\in\Big{[}0,\frac{1}{l}\Big{]}.$ (2.23)
It is easy to confirm the formula (2.22) numerically. We show the examples in Fig. 3.
Figure 3: The many-body entanglement entropy in the states with partial occupation (symbols) and the result from the few-body entanglement entropy (blue dashed lines). On the left panel we use the unit pattern $l=2,\kappa=\{0\}$, and on the right panel we use the unit pattern $l=3,\kappa=\{0,1\}$. Note the definition of the repetition ratio $z$ in (2.23).
The results indicate the effective form of the RDM
$\rho_{L,L_{A},p\kappa}\sim\underbrace{\rho_{l,\alpha+1,\kappa}\otimes\cdots\otimes\rho_{l,\alpha+1,\kappa}}_{\lfloor za\rfloor}\otimes\underbrace{\rho_{l,\alpha,\kappa}\otimes\cdots\otimes\rho_{l,\alpha,\kappa}}_{\lfloor z(\lfloor\frac{L}{l}\rfloor-a)\rfloor}\otimes\rho^{\prime}_{\lceil(1-z)L_{A}\rceil},$ (2.24)
where $\alpha\equiv\lfloor xl\rfloor$$a\equiv(L_{A}{\rm~mod~}\lfloor\frac{L}{l}\rfloor)$, and $\rho^{\prime}_{\lceil(1-z)L_{A}\rceil}$ is a density matrix of about $\lceil(1-z)L_{A}\rceil$ effective sites with sub-volume-law entanglement entropy.
2.3 States with mixed occupancy